OpenStudy (anonymous):

Given that the pair of triangles is similar, find the length of the side labeled x. 1st triangle is 33& x 2nd triangle is 27 & 36

OpenStudy (anonymous):

assuming these are corrsponding sides: 33 / 27 = x /36 x = 33* 36 / 27 = 33 * 4 / 3 = 44 ;

OpenStudy (anonymous):

x = 36 * 33/ 27 = 4 * 11 = 44